OpenStudy (anonymous):

A triangle has one side that measures x, and the other two sides each measure 4 inches less than x. The perimeter is 19 inches. What is the measure of x? 2 3 5 9

OpenStudy (anonymous):

@misty1212 @= @Luigi0210 @misssunshinexxoxo tagging random people

OpenStudy (anonymous):

@AbdullahM

OpenStudy (anonymous):

@MICAAAAAAAAAAAAAAAAAA

OpenStudy (anonymous):

This is a hard one for me.

OpenStudy (anonymous):

Sorry I can't help you!

OpenStudy (anonymous):

@pooja195 Help her please.

OpenStudy (abdullahm):

Hi!

OpenStudy (anonymous):

Hey

OpenStudy (abdullahm):

Do you know what perimeter means?

OpenStudy (anonymous):

Not really.

OpenStudy (anonymous):

isn't it like the base.

OpenStudy (misssunshinexxoxo):

Which answer do you believe it is?

OpenStudy (misssunshinexxoxo):

x+(x−4)+(x−4)=19

OpenStudy (abdullahm):

The perimeter means you add up all the three sides.

OpenStudy (anonymous):

@misssunshinexxoxo To be honest i'm not sure. I don't know how to work it out.

OpenStudy (abdullahm):

So, what is `4 inches less than x` mean? (In like an equation)

OpenStudy (anonymous):

\[4<x\]

OpenStudy (misssunshinexxoxo):

Yes

OpenStudy (anonymous):

okay. so what next

OpenStudy (abdullahm):

No, that means `x-4` 4 inches less than x means we have to do subtraction.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so 19-4?

OpenStudy (abdullahm):

So two of this sides are \(\sf x-4\) and one side is x So that means since the perimeter equals to 19: \(\sf\Large (x-4) + (x-4) + x=19\)

OpenStudy (anonymous):

So it would be 9?

OpenStudy (anonymous):

@AbdullahM how do i post a different question?

OpenStudy (abdullahm):

You have to first close this question and open another one.

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Thank you for helping me.

OpenStudy (abdullahm):

Also, show me your work for that last equation...

OpenStudy (anonymous):

okay

OpenStudy (abdullahm):

\(\sf\Large (x-4) + (x-4) + x=19\) we can make that: \(\sf\Large x-4 + x - 4 + x = 19\) Add like terms and then isolate x.

OpenStudy (anonymous):

(x+x+x)+(−4+−4)=19 there are 3 x's. -4+-4=-8 3x+−8=19 3x−8=19 then I added 8 to both sides. 3x=27 then I divide both sides by 3 x=9

OpenStudy (abdullahm):

Great job! :D

OpenStudy (anonymous):

:)